# Asynchronous Federated Stochastic Optimization with Exact Averaging for Heterogeneous Local Objectives

### v1

@@ 5 Numerical Results
In this section we evaluate the empirical performance of AREA. For our simulations we implemented a practical variant of AREA where the aggregation criterion is satisfied if $\Delta\in\mathbb{N}$ client updates have been received (see the discussion in Section 4.1 on how our convergence analysis can be interpreted as an upper bound for this aggregation policy). The elapsed time between two consecutive updates of client $i\in[n]$ was modeled as an exponential random variable $\text{exp}(\lambda_{i})$. We opted to compare AREA with the following methods: (i) synchronous FedAvg (S-FedAvg) [28], which we use as a synchronous benchmark, (ii) asynchronous FedAvg (AS-FedAvg) [e.g., 19] and (iii) FedBuff [33], two asynchronous methods sensitive to client drift, and (iv) MIFA [15], an asynchronous method robust to client drift. We note that in every figure in this section, the originally noisy lines have been smoothed to enhance visual clarity while maintaining the underlying data trends.
5.1 Quadratic problem
@@
5.2 Multinomial logistic regression
The corresponding optimization can be formulated as
$\min_{w^{1},\ldots,w^{C}\in\mathbb{R}^{d}}\quad f(w^{1},\ldots,w^{C})=\sum_{i=1}^{n}\frac{s_{i}}{S}\hat{f}_{i}(w^{1},\ldots,w^{C}),$
where $S=\sum_{i=1}^{n}s_{i}$.
We report results for the MNIST dataset [11] and network sizes $n=128$ and $n=4$ (cross-device vs. cross-silo settings [17]). For network size $n=4$ (resp. $n=128$) we set $\Delta=2$ (resp. $\Delta=4$) for AREA, AS-FedAvg, FedBuff, and MIFA. We set $\Delta=4$ for S-FedAvg in all cases. We used samples from a Dirichlet distribution with parameter $a=0.1$ to simulate non-iid local datasets (see Appendix A.4 for our results with iid data), and tested two different cases of client update frequencies: ($i$) uniform updates, where we set $\lambda_{i}=10$ for all $i\in[n]$, and ($ii$) non-uniform updates, where $\lambda_{i}$ for $i\in[n]$ were drawn from $\mathcal{N}(10,5)$. To evaluate performance, we tracked two metrics over (continuous) time: ($i$) the loss function value at the server model using all training data and ($ii$) the test accuracy of the server model using all test data. The batch size for the local stochastic gradients was set to $b=32$, and the regularization parameter in Problem (6) was set to $\nu=10^{-3}$. For all the methods, we tuned the stepsize $\alpha$ using grid search in the set $\{10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},10^{4}\}$ and selected the values resulting in the fastest convergence. We repeated the experiments $10$ times in total and extracted the minimum, maximum, and mean values of both selected metrics for the plots.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
5.2.1 Cross-device FL
Our results for the execution of a single local SGD step at each client and network size $n=128$ are shown in Fig. 3, respectively. The main takeaway from Fig. 3 is that AREA is able to tolerate larger stepsizes and achieve faster and more accurate convergence compared with the other methods for both uniform and non-uniform client update frequencies [^4] We note that the choice of stepsize for FedBuff ($\alpha=10^{1}$) resulted in instabilities at the beginning of the optimization, as seen in Figs. 3(a) and 3(c); however, the next smallest option ($\alpha=10^{0}$) results in significantly slower convergence.. On the other hand, our results indicate that MIFA requires small stepsizes to guarantee numerical stability. An additional observation is that the performance of S-FedAvg degrades substantially when the client update frequencies are non-uniform. Finally, although providing theoretical guarantees for AREA in the case where multiple local SGD steps are performed at each client is out of the scope of this work, we tested AREA and the remaining methods with $50$ local SGD steps at each client (Fig. 4). We note that performing multiple SGD steps alleviates communication load but introduces client drift to AREA. In Fig. 4 we observe that as in the single local SGD step case, AREA is able to tolerate larger stepsizes relative to the other methods and achieves faster convergence with respect to testing accuracy (Figs.  4(b) and 4(d)). However, the large stepsize and the introduction of client drift also result in a large error neighborhood in the loss function plots (Figs. 4(a) and 4(c)).
5.2.2 Cross-silo FL
@@ 6 Conclusion
In this paper we proposed and analyzed AREA, a new asynchronous stochastic (sub)gradient method for FL that is robust to client drift and guaranteed to converge under arbitrarily long client delays. For smooth and strongly convex functions, we established the linear convergence rate of AREA to an error neighborhood of the optimal solution of Problem (1), whose size depends only on the variance of the stochastic gradient approximations used and is independent of client delays and the heterogeneity between the local datasets. For constrained problems with convex and nonsmooth functions, we recovered the same convergence rate as the centralized projected subgradient method. Our numerical results on a quadratic and a multinomial logistic regression problem suggest that AREA performs favorably compared with state-of-the-art methods, particularly when the local datasets are highly non-iid.
@@ Appendix A Appendix
A.1 Preliminaries
@@
A.2 Main results for strongly convex & smooth functions
@@
$\begin{split}\|x_{i,k+1}-x_{i}^{\star}\|^{2}&=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha g_{i}(x_{s,k})-x_{i}^{\star}\|^{2}\\ &=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star}\|^{2}+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle\\ &\leq(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}(1-\alpha\gamma_{i})\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle,\end{split}$
@@
A.3 Main results for convex problems
@@
$\begin{split}f_{j}(y_{i,k})-f_{j}(x_{s,k})&\leq p_{i}^{-1}(f_{j}(y_{i,k})+f_{j}(x_{i,k})-2f_{j}^{\star})-p_{i}^{-1}\mathbb{E}[f_{j}(y_{i,k+1})+f_{j}(x_{i,k+1})-2f_{j}^{\star}|\mathcal{F}_{k}]+\alpha_{k}\sigma_{1}^{2},\end{split}$ (12b)
@@
$\begin{split}f_{j}(x_{i,k+1})&=(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}f_{j}(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})])\\ &\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\left(f_{j}(x_{s,k})-\langle d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right),x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\rangle\right),\end{split}$
@@
$\begin{split}f_{j}(x_{i,k+1})&\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\left(f_{j}(x_{s,k})+\left\|d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right)\right\|\|x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\|\right)\\ &\leq f_{j}(x_{i,k})+\phi_{i,k+1}(f_{j}(x_{s,k})-f_{j}(x_{i,k}))+\alpha_{k}\phi_{i,k+1}C(C+\|g_{i}(x_{s,k})-d_{i}(x_{s,k})\|),\end{split}$
@@
$\begin{split}\|x_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}(\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle)\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}(\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle)\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2},\end{split}$
@@
$\begin{split}\|y_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|y_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|y_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2},\end{split}$
@@
$\begin{split}&\mathbb{E}[\|x_{i,k+1}-y_{i,k+1}\|^{2}+\alpha_{k+1}B^{x}_{ii,k+1}|\mathcal{F}_{k}]\leq p_{i}\|x_{i,k}-x_{s,k}\|^{2}+(1-p_{i})\|x_{i,k}-y_{i,k}\|^{2}+\alpha_{k}B^{x}_{ii,k+1}+\alpha_{k}^{2}p_{i}\sigma_{3}^{2}.\end{split}$ (15)
@@
$\begin{split}\|x_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-y_{j,k}\|^{2}+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2}\\ &\leq\phi_{i,k+1}\|y_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}\phi_{i,k+1}(f_{i}(y_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-y_{j,k},\epsilon_{i,k}\rangle+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2},\end{split}$
@@
$\begin{split}\|x_{i,k+1}-x_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{j,k}\|^{2}+\phi_{j,k+1}\|x_{i,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{j}(x_{s,k})]\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k+1}-x_{j,k+1}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{j,k},\epsilon_{i,k}\rangle\\ &\quad+\phi_{j,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{j,k+1}(C^{2}+\|\epsilon_{j,k}\|^{2})+2\alpha_{k}\phi_{j,k+1}(f_{j}(x_{i,k})-f_{j}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{j,k+1}\langle x_{s,k}-\alpha_{k}d_{j}(x_{s,k})-x_{i,k},\epsilon_{j,k}\rangle+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-x_{j,k}\|^{2},\end{split}$
@@
$\begin{split}\|y_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{j,k+1}\|y_{i,k}-x_{j,k}\|^{2}+(1-\phi_{i,k+1}-\phi_{j,k+1})\|y_{i,k}-y_{j,k}\|^{2}.\end{split}$
@@
$\begin{split}\|x_{i,k+1}-y_{i,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\|^{2}+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{i,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{i,k},\epsilon_{i,k}\rangle+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2},\end{split}$
@@
$\begin{split}\mathbb{E}[\|x_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{i})\|x_{i,k}-x^{\star}\|^{2}+p_{i}\|x_{s,k}-x^{\star}\|^{2}+p_{i}\alpha^{2}_{k}\hat{\sigma}^{2}+2p_{i}\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k})),\end{split}$ (17)

### v2

@@ 5 Numerical Results
In this section we evaluate the empirical performance of AREA. For our simulations we implemented a practical variant of AREA where the aggregation criterion is satisfied if $\Delta\in\mathbb{N}$ client updates have been received (see the discussion in Section 4 on how our convergence analysis can be interpreted as an upper bound for this aggregation policy). The elapsed time between two consecutive updates of client $i\in[n]$ was modeled as an exponential random variable $\text{exp}(\lambda_{i})$. We opted to compare AREA with the following methods: (i) synchronous FedAvg (S-FedAvg) [29], which we use as a synchronous benchmark, (ii) asynchronous FedAvg (AS-FedAvg) (e.g., [20]) and (iii) FedBuff [35], two asynchronous methods sensitive to client drift, and (iv) MIFA [17], an asynchronous method robust to client drift. We note that in every figure in this section, the originally noisy lines have been smoothed to enhance visual clarity while maintaining the underlying data trends.
5.1 Multinomial logistic regression
The corresponding optimization can be formulated as $\min_{w^{1},\ldots,w^{C}\in\mathbb{R}^{d}}\quad f(w^{1},\ldots,w^{C})=\sum_{i=1}^{n}\frac{s_{i}}{S}\hat{f}_{i}(w^{1},\ldots,w^{C}),$ where $S=\sum_{i=1}^{n}s_{i}$. We report results for the MNIST dataset [12] and network size $n=128$. We set $\Delta=4$ for all methods, and used samples from a Dirichlet distribution with parameter $a=0.1$ to simulate non-iid local datasets. We tested two different cases of client update frequencies: ($i$) uniform updates, where we set $\lambda_{i}=10$ for all $i\in[n]$, and ($ii$) non-uniform updates, where $\lambda_{i}$ for $i\in[n]$ were drawn from $\mathcal{N}(10,5)$. To evaluate performance, we tracked two metrics over (continuous) time: ($i$) the loss function value at the server model using all training data and ($ii$) the test accuracy of the server model using all test data. The batch size for the local stochastic gradients was set to $b=32$, and the regularization parameter in Problem (8) was set to $\nu=10^{-3}$. For all the methods, we tuned the stepsize $\alpha$ using grid search in the set $\{10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},10^{4}\}$ and selected the values resulting in the fastest convergence. We repeated each experiment configuration $10$ times on a CPU cluster with $36$ compute nodes of memory 128GB DDR4 ($\sim 4$ hours per configuration, $\sim 4$ days total computation for the experiments in this section) and extracted the minimum, maximum, and mean values of both selected metrics for the plots. Additional results for non-iid data and $n=4$ can be found in Appendix A.2.2, and our results for iid data can be found in Appendix A.3.
@@ 5.1 Multinomial logistic regression
Our results for the execution of a single local SGD step at each client are shown in Fig. 1. The main takeaway from Fig. 1 is that AREA is able to tolerate larger stepsizes and achieve faster and more accurate convergence compared with the other methods for both uniform and non-uniform client update frequencies [^3] We note that the choice of stepsize for FedBuff ($\alpha=10^{1}$) resulted in instabilities at the beginning of the optimization, as seen in Figs. 1(a) and 1(c); however, the next smallest option ($\alpha=10^{0}$) results in significantly slower convergence.. Our results also indicate that MIFA requires small stepsizes to guarantee numerical stability. An additional observation is that the performance of S-FedAvg degrades substantially when the client update frequencies are non-uniform. Finally, although providing theoretical guarantees for AREA in the case where multiple local SGD steps are performed at each client is out of the scope of this work, we tested AREA and the remaining methods with $50$ local SGD steps at each client (Fig. 2). We note that performing multiple SGD steps alleviates communication load but introduces client drift to AREA. In Fig. 2 we observe that as in the single local SGD step case, AREA is able to tolerate larger stepsizes relative to the other methods and achieves faster convergence with respect to testing accuracy (Figs.  2(b) and 2(d)). However, the large stepsize and the introduction of client drift also result in a large error neighborhood in the loss function plots (Figs. 2(a) and 2(c)).
6 Limitations
A significant limitation of our work is that we do not provide theoretical or numerical results for nonconvex functions (e.g. neural networks), which is a more practical setup for FL applications. The study of the convergence properties of AREA in the nonconvex setting is left for future work. Moreover, due to the probabilistic nature of our analysis, which allows for arbitrarily long delays between iterations, it is not possible to readily obtain time complexities unless additional assumptions are adopted on the distributions of client delays, e.g., Example 4.
@@ 7 Conclusion
We proposed and analyzed AREA, a new asynchronous stochastic (sub)gradient method that is robust to the client drift caused by heterogeneous client update frequencies and converges under arbitrary delays. For smooth and strongly convex functions, we established the asymptotic convergence of AREA to an error neighborhood whose size depends only on the variance of the stochastic gradient approximations used and is independent of the maximum client delay and the heterogeneity between the local datasets. For constrained problems with convex and non-smooth functions, we recovered the same convergence rate as the centralized projected subgradient method. Our numerical results on a multinomial logistic regression problem suggest that AREA performs favorably compared with state-of-the-art methods, particularly when the local datasets are highly non-iid.
@@ Appendix A Appendix
A.1 Proofs and additional theoretical results
A.1.1 Preliminaries
@@
A.1.2 Main results for strongly convex & smooth functions
@@
$\begin{split}\|x_{i,k+1}-x_{i}^{\star}\|^{2}&=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha g_{i}(x_{s,k})-x_{i}^{\star}\|^{2}\\ &=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star}\|^{2}\\ &\quad+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle\\ &\leq(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}(1-\alpha\gamma_{i})\|x_{s,k}-x^{\star}\|^{2}\\ &\quad+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle,\end{split}$
@@
A.1.3 Main results for convex problems
@@
$\begin{split}f_{j}(y_{i,k})-f_{j}(x_{s,k})&\leq p_{i}^{-1}(f_{j}(y_{i,k})+f_{j}(x_{i,k})-2f_{j}^{\star})\\ &\quad-p_{i}^{-1}\mathbb{E}[f_{j}(y_{i,k+1})+f_{j}(x_{i,k+1})-2f_{j}^{\star}|\mathcal{F}_{k}]+\alpha_{k}\sigma_{1}^{2},\end{split}$ (14b)
@@
$\begin{split}f_{j}(x_{i,k+1})&=(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}f_{j}(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})])\\ &\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\bigg{(}f_{j}(x_{s,k})\\ &\quad-\langle d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right),x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\rangle\bigg{)},\end{split}$
@@
$\begin{split}f_{j}(x_{i,k+1})&\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\bigg{(}f_{j}(x_{s,k})\\ &\quad+\left\|d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right)\right\|\|x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\|\bigg{)}\\ &\leq f_{j}(x_{i,k})+\phi_{i,k+1}(f_{j}(x_{s,k})-f_{j}(x_{i,k}))+\alpha_{k}\phi_{i,k+1}C(C+\|g_{i}(x_{s,k})-d_{i}(x_{s,k})\|),\end{split}$
@@
$\begin{split}\|x_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}(\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle)\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}\bigg{(}\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}\\ &\quad+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle\bigg{)}+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2},\end{split}$
@@
$\begin{split}\|y_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|y_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|y_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2},\end{split}$
@@
$\begin{split}\mathbb{E}[\|x_{i,k+1}-y_{i,k+1}\|^{2}+\alpha_{k+1}B^{x}_{ii,k+1}|\mathcal{F}_{k}]&\leq p_{i}\|x_{i,k}-x_{s,k}\|^{2}+(1-p_{i})\|x_{i,k}-y_{i,k}\|^{2}\\ \ &\quad+\alpha_{k}B^{x}_{ii,k+1}+\alpha_{k}^{2}p_{i}\sigma_{3}^{2}.\end{split}$ (17)
@@
$\begin{split}\|x_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-y_{j,k}\|^{2}+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2}\\ &\leq\phi_{i,k+1}\|y_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(y_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-y_{j,k},\epsilon_{i,k}\rangle\\ &\quad+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2},\end{split}$
@@
$\begin{split}\|x_{i,k+1}-x_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{j,k}\|^{2}\\ &\quad+\phi_{j,k+1}\|x_{i,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{j}(x_{s,k})]\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k+1}-x_{j,k+1}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{j,k},\epsilon_{i,k}\rangle\\ &\quad+\phi_{j,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{j,k+1}(C^{2}+\|\epsilon_{j,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{j,k+1}(f_{j}(x_{i,k})-f_{j}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{j,k+1}\langle x_{s,k}-\alpha_{k}d_{j}(x_{s,k})-x_{i,k},\epsilon_{j,k}\rangle\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-x_{j,k}\|^{2},\end{split}$
@@
$\begin{split}\|y_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{j,k+1}\|y_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|y_{i,k}-y_{j,k}\|^{2}.\end{split}$
@@
$\begin{split}\|x_{i,k+1}-y_{i,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\|^{2}+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{i,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{i,k},\epsilon_{i,k}\rangle+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2},\end{split}$
@@
$\begin{split}\mathbb{E}[\|x_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{i})\|x_{i,k}-x^{\star}\|^{2}+p_{i}\|x_{s,k}-x^{\star}\|^{2}\\ &\quad+p_{i}\alpha^{2}_{k}\hat{\sigma}^{2}+2p_{i}\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k})),\end{split}$ (19)